# Application of Langevin Dynamics to Advance the Quantum Natural Gradient Optimization Algorithm

### v3

@@ 3.2 Quantum Approximate Optimization Algorithm
Our second test model is the Minimum Vertex Cover problem treated in the framework of the QAOA approach. Recently, this problem has been used to study the impact of noise on classical optimizers and to determine the optimal depth of the QAOA circuit [25]. In our calculations we use a modified code by Jack Ceroni [26] to study two graphs with $N=4$ and $N=8$ vertices. We build QAOA circuits with 4 layers for $N=4$ and with 6 layers for $N=8$. For this problem we run a series of 200 trials with the same for all optimizers random-guessed initial values of variational parameters. The optimization process runs for 200($N=4$)/300($N=8$) steps or until energy convergence up to the 2-digit ($N=4$) / 3-digit ($N=8$) accuracy during at least 3 steps. To compare the performance of different optimizers, we calculate the quality ratio of the final optimized state - the total probability to find the states of the exact solution in the given optimized solution. The values of hyperparameters are as follows: $0.01\leq\eta\leq 0.25$, $\rho=0.9$ for Momentum and Momentum-QNG, and $\beta_{1}=0.9$, $\beta_{2}=0.99$, $\epsilon=10^{-8}$ for Adam.
From Fig. 2(a) for $N=4$ one can see that in the range of stepsize (learning rate $\eta$) values under consideration Adam, Momentum and Momentum-QNG achieve almost equal maximal values of the quality ratio. At the same time we note that, for $\eta<0.2$, Momentum-QNG systematically outperforms both Adam and original QNG. From Fig. 2(c) for $N=8$ one can see that Momentum goes to maximum and then decays very fast. Again, Momentum-QNG outperforms the basic QNG for $\eta<0.2$ and performs similar to Adam for $\eta\leq 0.15$.
To study the convergence behaviour of the optimization algorithms under consideration, on Fig. 2(b) and (d) we plot the mean (symbols) and the standard deviation (error bars) of the number of steps to convergence of the optimization procedure in a series of 200 trials, as a function of the value of stepsize (learning rate, $\eta$). For both $N=4$ and $N=8$, with the increase of stepsize, Momentum quickly goes above the convergence limit (200 and 300 steps, respectively), QNG and Momentum-QNG demonstrate the intermediate results, while Adam shows the best convergence behaviour. For further details of our calculations see [27].
Figure 2: Benchmarking Momentum-QNG together with QNG, Momentum and Adam on the Minimum Vertex Cover problem. The vertical axis shows the mean (symbols) and the standard deviation (error bars) of the quality ratio (a) ($N=4$), (c) ($N=8$) and of the number of steps to convergence (b) ($N=4$), (d) ($N=8$) in a series of 200 trials, while the horizontal axis shows the value of stepsize (learning rate, $\eta$) of four different optimizers under consideration. Each trial is initialized with random-guess values of the variational parameters, being the same for all optimizers. Then the optimization process runs for 200 (b) or 300 (d) steps (or until convergence) independently by each optimizer.
@@ 4 Conclusions
In this paper we demonstrate that application of Langevin dynamics with Quantum Natural Gradient force for optimization of variational quantum circuits gives a new optimization algorithm, which we call Momentum-QNG. The basic QNG algorithm uses the quantum metric tensor to rescale the variational parameter space to give a more symmetric shape of the loss function. On the other hand, the momentum (inertial) term in the Momentum algorithm prevents the optimization process from sticking to the local minima and plateaus. Here we report that the combination of both these features in the Momentum-QNG algorithm gives a strong boost effect. Indeed, for the Minimum Vertex Cover problem (see Fig. 2(a) and (c)) the proposed Momentum-QNG algorithm outperforms the basic QNG in the range of the learning rate values $0.2250.01\leq\eta<$ 0.225. And for the Investment Portfolio Optimization problem (see Fig. 1(a) and (c)) it gives the best result between all the optimizers under consideration for $0.01\leq\eta\leq 0.25$ ($N=6$) and $0.01\leq\eta\leq 0.1$ ($N=12$).
Based on our results, we propose $\rho=0.9$ and $\eta=0.1$ as default hyperparameter values for Momentum-QNG.
Author contributions
OB proposed the concept of Momentum-QNG. MB, IL, ML and IO developed the code and did benchmarking calculations. AS and AL proposed benchmarking applications. All authors analyzed and discussed the results obtained. All authors read and approved the final manuscript.
Acknowledgement
@@
This study was initiated during the QHACK bootcamp organized by Haiqu and continued during QHACK2024 organized by Xanadu. We deeply appreciate the inspiring atmosphere of both these events and further seminars and discussions. O.B., M.B. and I.O. have received funding through the EURIZON project, which is funded by the European Union under grant agreement No.871072. I.L. and A.S. acknowledge support by the National Research Foundation of Ukraine, project No. 2023.03/0073.
Appendix A Cumulative probability distributions
In this appendix section we present the cumulative probability distributions calculated for different stepsize values.
A.1 Investment Portfolio Optimization problem
From Fig. 3 and Fig. 4 one can see that Momentum-QNG gives the highest probability to find the energy of the final optimized state within the range $\Delta E\leq 0.2$. At the same time from Fig. 5 and Fig. 6 one can see that Momentum-QNG demonstrates the least variation of the number of steps to convergence in the whole stepsize range (see also Fig. 1(b)).
A.2 Minimum Vertex Cover problem
From Fig. 7 and Fig. 8 one can see that Momentum-QNG demonstrates one of the best quality ratio distributions among the algorithms studied and definitely outperforms the basic QNG optimizer. At the same time, from Fig. 9 and Fig. 10 one can see that Momentum-QNG demonstrates a modest convergence behaviour compared to Adam, similar to the basic QNG and Momentum.
Figure 3: Benchmarking Momentum-QNG (c) together with QNG (a), Momentum (d) and Adam (b) on the Portfolio Optimization problem with $N=6$. The horizontal axis shows the value of the difference between the optimized and the ground state energy, while the vertical axis shows the cumulative probability to obtain the corresponding energy difference in a series of 200 trials. The curves are calculated for all values of the stepsize (learning rate, $\eta$) under consideration.
Figure 4: Benchmarking Momentum-QNG (c) together with QNG (a), Momentum (d) and Adam (b) on the Portfolio Optimization problem with $N=12$. The horizontal axis shows the value of the difference between the optimized and the ground state energy, while the vertical axis shows the cumulative probability to obtain the corresponding energy difference in a series of 200 trials. The curves are calculated for all values of the stepsize (learning rate, $\eta$) under consideration.
Figure 5: Benchmarking Momentum-QNG (c) together with QNG (a), Momentum (d) and Adam (b) on the Portfolio Optimization problem with $N=6$. The horizontal axis shows the number of steps to energy convergence up to the 3-digit accuracy, while the vertical axis shows the cumulative probability to obtain the corresponding number of steps in a series of 200 trials. The curves are calculated for all values of the stepsize (learning rate, $\eta$) under consideration.
Figure 6: Benchmarking Momentum-QNG (c) together with QNG (a), Momentum (d) and Adam (b) on the Portfolio Optimization problem with $N=12$. The horizontal axis shows the number of steps to energy convergence up to the 3-digit accuracy, while the vertical axis shows the cumulative probability to obtain the corresponding number of steps in a series of 200 trials. The curves are calculated for all values of the stepsize (learning rate, $\eta$) under consideration.
Figure 7: Benchmarking Momentum-QNG (c) together with QNG (a), Momentum (d) and Adam (b) on the Minimum Vertex Cover problem for $N=4$. The horizontal axis shows the value of the quality ratio, being a characteristic of an optimized state, while the vertical axis shows the cumulative probability to obtain the corresponding quality ratio in a series of 200 trials. The curves are calculated for all values of the stepsize (learning rate, $\eta$) under consideration.
Figure 8: Benchmarking Momentum-QNG (c) together with QNG (a), Momentum (d) and Adam (b) on the Minimum Vertex Cover problem for $N=8$. The horizontal axis shows the value of the quality ratio, being a characteristic of an optimized state, while the vertical axis shows the cumulative probability to obtain the corresponding quality ratio in a series of 200 trials. The curves are calculated for all values of the stepsize (learning rate, $\eta$) under consideration.
Figure 9: Benchmarking Momentum-QNG (c) together with QNG (a), Momentum (d) and Adam (b) on the Minimum Vertex Cover problem for $N=4$. The horizontal axis shows the number of steps to energy convergence up to the 2-digit accuracy during at least 3 steps, while the vertical axis shows the cumulative probability to obtain the corresponding number of steps in a series of 200 trials. The curves are calculated for all values of the stepsize (learning rate, $\eta$) under consideration.
Figure 10: Benchmarking Momentum-QNG (c) together with QNG (a), Momentum (d) and Adam (b) on the Minimum Vertex Cover problem for $N=8$. The horizontal axis shows the number of steps to energy convergence up to the 3-digit accuracy during at least 3 steps, while the vertical axis shows the cumulative probability to obtain the corresponding number of steps in a series of 200 trials. The curves are calculated for all values of the stepsize (learning rate, $\eta$) under consideration.

### v4

@@ 3.2 Quantum Approximate Optimization Algorithm
Our next test model is the Minimum Vertex Cover problem treated in the framework of the QAOA approach. Recently, this problem has been used to study the impact of noise on classical optimizers and to determine the optimal depth of the QAOA circuit [37]. In our calculations we use a modified code by Jack Ceroni [38] to study two graphs with $N=4$ and $N=8$ vertices. We build QAOA circuits with 4 layers for $N=4$ qubits and with 6 layers for $N=8$ qubits. Then we run a series of 200 trials with the same for all optimizers random-guessed initial values of variational parameters. The optimization process runs for 200 steps or until energy convergence up to the 2-digit accuracy during at least 3 steps. To compare the performance of different optimizers, we calculate the quality ratio of the final optimized state – the total probability to find the states of the exact solution in the given optimized solution. The range of learning rate values studied is $0.01\leq\eta\leq 2$ for $N=4$ and $0.001\leq\eta\leq 0.9$ for $N=8$.
From Fig. 3(a) for $N=4$ one can see that Adam, Momentum and Momentum-QNG achieve almost equal maximal values of the quality ratio within their convergence domains and significantly outperform the momentumless QNG. From Fig. 3(c) for $N=8$ one can see that Momentum-QNG performs very similar to Adam within its convergence domain and achieves almost the same maximal quality ratio as Momentum. At the same time, Adam achieves the highest quality ratio within its convergence domain. Again, the momentumless QNG achieves the least quality ratio.
To study the convergence behaviour of the optimization algorithms under consideration, in Fig. 3(b) and (d) we plot the mean (symbols) and the standard deviation (shaded regions) of the number of steps to convergence, as a function of the learning rate $\eta$. For both $N=4$ and $N=8$, Momentum demonstrates the narrowest convergence domain, QNG and Momentum-QNG exhibit similar intermediate-range convergence domains, while Adam shows the widest convergence domain. For further details of our calculations see our code [39].
Figure 3: Benchmarking Momentum-QNG together with QNG, Momentum and Adam on the Minimum Vertex Cover problem. The vertical axis shows the mean (symbols) and the standard deviation (shaded regions) of the quality ratio (a) ($N=4$), (c) ($N=8$) and of the number of steps to convergence (b) ($N=4$), (d) ($N=8$) in a series of 200 trials, while the horizontal axis shows the learning rate $\eta$ of four different optimizers under consideration.
@@ 4 Conclusions
In this paper we demonstrate that application of Langevin dynamics with Quantum Natural Gradient force for optimization of variational quantum circuits gives a new optimization algorithm, which we call Momentum-QNG.
The basic QNG algorithm uses the Quantum Geometric tensor to rescale the variational parameter space to give a more symmetric shape of the objective function. On the other hand, the momentum (inertial) term in the Momentum algorithm prevents the optimization process from sticking to the local minima and plateaus. Indeed, from Eq. (17) one can see that the mean jump length in the vicinity of a local extremum of the objective function increases with increasing momentum coefficient. This feature allows momentum-amended algorithms to explore a wider volume in the variational parameter space and to find deeper minima.
This conclusion is supported by our numerical experiments. Indeed, for both the Investment Portfolio (see Fig. 1(a), (c) and (e)) and the Minimum Vertex Cover (see Fig. 3(a) and (c)) optimization problems the momentum-amended algorithms (Adam, Momentum and Momentum-QNG) outperform the momentumless QNG. It is worth noting that Adam demonstrates the best performance in both these problems. For the quantum Sherrington-Kirkpatrick model with relatively strong transverse field $g=0.1$, the basic QNG algorithm demonstrates the best optimization performance (Fig. 2(a)), in agreement with Stokes et al. [12]. For $g=10^{-3}$ the Momentum and Momentum-QNG algorithms show almost equal best results (Fig. 2(c)) and for $g=10^{-5}$ Momentum-QNG outperforms the rest (Fig. 2(e)). One should take into account that at small values of the transverse field the spin-glass features of the quantum Sherrington-Kirkpatrick model, including multiple local minima of the energy landscape, become more pronounced. It is tempting to assume that a synergetic effect of application of the quantum geometric tensor and momentum results into this enhanced performance.
Acknowledgements
@@
This study was initiated during the QHACK bootcamp organized by HAIQU and continued during QHACK2024 organized by Xanadu Quantum Technologies. We deeply appreciate the inspiring atmosphere of both these events and further seminars and discussions. We are grateful to Dr. Mykyta Bulakhov for his help with multiprocessor cluster calculations. O.B., M.B. and I.O. acknowledge funding through the EURIZON project, which is funded by the European Union under grant agreement No.871072. I.L. acknowledges support by the IMPRESS-U grant from the US National Academy of Sciences and Office of Naval Research Global via STCU project No. 7120 and the IEEE program ‘Magnetism for Ukraine 2025’, Grant No. 9918. A.S. acknowledges support by the National Research Foundation of Ukraine, project No. 0124U004372.
Author contributions statement
Oleksandr Borysenko: Writing – original draft, Methodology, Conceptualization, Investigation, Funding acquisition, Project administration. Mykhailo Bratchenko: Writing – original draft, Investigation, Methodology, Software, Visualization. Ilya Lukin: Writing – original draft, Investigation, Methodology, Software, Visualization. Mykola Luhanko: Writing – original draft, Investigation, Software, Visualization. Ihor Omelchenko: Writing – original draft, Investigation, Software, Visualization. Andrii Sotnikov: Writing – original draft, Investigation, Funding acquisition, Supervision, Project administration. Alessandro Lomi: Writing – original draft, Investigation, Supervision.
Declaration of competing interest
The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.
Data availability
Our open source code and more results are available at https://github.com/borbysh/Momentum-QNG